OpenStudy (anonymous):

Solve -2(3-m)+15=6m-4(m-20) Please show all of your work

OpenStudy (anonymous):

-2(3-m)+15=6m-4(m-20) -6+2m+15=6m-4m+80 2m+9=2m+80 No solution.

OpenStudy (anonymous):

Okay thanks! I thought so I just wanted to make sure :)